[Brahms-l] Shift report 20040213 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Feb 13 2004 - 23:45:27 EST
------------------------------------------------------------
Date: Fri Feb 13 23:45:27 2004, Shift: 16:00-24:00
Supervisor: ZIS, with: EJK
Log book#15 pages 56 - 56

Shift Summary:
------------------------------------------------------------
1. We had beam ~4.5 h

2. Data run 10325-10328 : FS4 tr2(233k), MRS 45 trg3(485k)
   Pedestal : 10322
   TPC pedestal : 10323

3. No Detector Problems
   PII 3 has been down!
   No clue for that.
   Flemming and Hiro fixed it temporary.
   SCSI cable was changed.
   (->please read FV's email)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 13 23:45:33 2004

This archive was generated by hypermail 2.1.8 : Fri Feb 13 2004 - 23:45:55 EST